OpenStudy (anonymous):

Find the x- and y- intercepts for the graph of the following equation. 5x + 3y = -30

OpenStudy (anonymous):

need help

jimthompson5910 (jim_thompson5910):

To find the x-intercept, plug in y = 0 and solve for x 5x + 3y = -30 5x + 3(0) = -30 5x + 0 = -30 5x = -30 x = -30/5 x = -6 So the x-intercept is x = -6, which is the point (-6, 0)

jimthompson5910 (jim_thompson5910):

To find the y-intercept, plug in x = 0 and solve for y. I'll let you do this part.

OpenStudy (anonymous):

so then y=0

jimthompson5910 (jim_thompson5910):

y = 0 for any x-intercept, yes

jimthompson5910 (jim_thompson5910):

x = 0 for any y intercept

OpenStudy (anonymous):

x=-6 y=0 correct

jimthompson5910 (jim_thompson5910):

that's one ordered pair

jimthompson5910 (jim_thompson5910):

that's the x-intercept

OpenStudy (anonymous):

soo whats the answers ? this isn't homework by the way